OpenStudy (faman39):

Amount of interest: a. $1, 800 at 2% for 10 years . b.$1000 at 6% for 4 years. c. $2,500 at 3% for 5 years. d. $900 at 10 % for 3 years.

OpenStudy (anonymous):

What grade is this for?

OpenStudy (anonymous):

a) $2,160 b) $1,240 c) $2,875 d) $1,170 hope this helped

OpenStudy (anonymous):

. . . Simple interest: I = prt a. $1, 800 at 2% for 10 years = 18 * 2 * 10 = $360 b. $1000 at 6% for 4 years = 10 * 6 * 4 = $240 c. $2,500 at 3% for 5 years: = 25 * 3 * 5 = $375 d. $900 at 10 % for 3 years: = 9 * 10 * 3 = $ 270